# On the Quest for Effectiveness in Human Oversight: Interdisciplinary Perspectives

### v1

@@ 3.1.1. Causal power
Applied to the context of human oversight, this means that the human overseer has the power to establish a causal connection to parts of the system or its effects, especially the risk. For instance, this may include the ability to influence the system, including its components, operation, or outputs. Which kind of causal power is appropriate and necessary depends on the context at hand. For example, a factory robot assisting a human in assembly might need a stop button or a manual control mechanism. For an AI tool that assesses university applications and makes recommendations for the admission committee it might be enough if the human overseer can overwrite or disregard the system’s output and then forward possible issues to the system provider.
In any case, it is not enough if the human overseer is a mere observer of the system. Without proper means to interfere with it or at least with its effects, they could only watch the manifestation of a risk without being able to mitigate it. Also, it is not sufficient if the human only has ‘pseudo-options’, i.e., if they can only seemingly influence the system but not really make a relevant difference. [^3] Pseudo-options, though, are not to be confused with merely unattractive but real options, such as pushing a stop-button for fun and thereby halting an entire production line for no reason. Instances of such scenarios are malfunctioning stop buttons, or environments in which the overseer’s actions are by default not considered.
@@ 3.1.2. Epistemic access
To be effective, the human overseer must know what to do and when to do it. More precisely:
@@
Applied to the context of human oversight, this means that the human overseer has sufficient knowledge about the risk and how to mitigate it, especially knowing a risk mitigation action. A human overseer will, for instance, need to have sufficient understanding of what the system is doing, in which state it is, which possible risks and benefits there are, which means of influencing the system exist, which effects will result from possible interventions, and which of these effects are most desirable. In some cases, this might include knowledge of normative, social, or cultural domains, without which it would be infeasible for the overseer to correctly assess a certain risk. The overseer does not need to have perfect knowledge, though. They might be unsure or even ignorant about some aspects of their decision situation – as long as they know enough to achieve their relevant objectives, they have enough epistemic access. Notably, we assume that the overseer has certain basic knowledge, such as what a stop button is, or how to operate a computer on a consumer-level.
@@ 3.1.3. Self-control
To be effective, the human overseer must be in charge of their own doing. More precisely:
@@
Applied to the context of human oversight, this means that if here is a risk-mitigating path of action, the human can decide for that path, and if they do so, they actually perform the corresponding action(s). For this, the human overseer needs to be in their right state of mind and free in their actions. They also have to be in a position to retain attention and purposefully target their actions during the oversight. If, for example, the job of overseeing a fleet of autonomous vehicles is so uneventful and dull that a normal person cannot help but let their mind wander off from their task, the oversight would ultimately become ineffective. Only if human overseers have self-control, they can purposefully address risks. Other reasons why an overseer could lack self-control would be when they are severely sleep-deprived, under the influence of substances, or suffering from an active seizure.
An overseer who lacks self-control might also simultaneously lack sufficient causal power or epistemic access (e.g., in case of a seizure). Nevertheless, self-control is distinct from the other two conditions, as an overseer can have self-control without having causal power or epistemic access. In fact, self-control may, in many situations, be an antecedent of causal power and epistemic access – but not vice versa.
Sometimes, human overseers will decide in split-seconds or will instinctively perform a certain move, and we typically regard this as self-controlled. For example, a human overseer of a factory robot who realizes that it is just milliseconds away from severely injuring a human worker might reflexively slam the stop-button. We consider this to be a real decision, even if made in a split-second; and since the overseer followed suit with their decision, they are to be considered self-controlled. Therefore, even overseers deciding instinctively can be effective according to the upcoming Definition 3.5.555Alternatively, one might say there was not a true decision made in the example, but rather an instinctive jolt for the stop-button. In this case, the human overseer would not count as self-controlled according to Definition 3.3 (and therefore also not as effective according to the upcoming Definition 3.5). If one wishes to allow the second interpretation of the situation above and simultaneously allow that the human in it was exercising effective oversight, then the definition of self-control might be adapted to say “could” instead of “can”. In this case, even certain instinctive movements of the above kind fall under the definition. However, since we want to avoid a discussion of the conceptual idealization introduced by the “could”, we will stick to Definition 3.3 as is.
@@ 3.1.4. Fitting intentions
Applied to the context of human oversight, this means that the human overseer pro tanto intents to mitigate risks, i.e., they intent to mitigate any risks while taking other relevant factors, such as the interests of the system’s users and other stakeholders, into consideration in a suitable way. If the overseer was unwilling to take the efforts necessary to mitigate a risk, they would not be effective in their oversight – even if they had causal power, epistemic access, and self-control. This may, for example, be the case if the overseer lacks motivation to do their job properly, or has other, conflicting interests. Even worse, if a human overseer had ill intent, they could use their position to foster risks instead of mitigating them. For example, a racist overseeing a university admission system could deliberately overwrite outputs where people of color are ranked highly. In this case, their oversight should be considered ineffective in the relevant respects.
@@ 3.1.5. Effectiveness
A human overseer is effective in their human oversight if and only if they have causal power, epistemic access, self-control, and fitting intentions in the above senses.
@@
Moreover, “effectiveness” is not a binary concept but a matter of degrees: one exercise of oversight might be more effective than another. Therefore, considering effectiveness in human oversight as a multifaceted continuum suitably encapsulates its nuanced nature. In line with the literature (Green, 2022) and legislation (European Commission, 2021), we nevertheless allow the use of “effective” as a binary concept. Saying that they are effective is to mean that they are effective enough for the concrete context they are operating in. In essence, this is also what Definition 3.5 says. While it is an open and important question how much effectiveness is enough for a given context, we will not discuss this issue as it would go far beyond the scope of this paper.
@@ 3.2.1. Competence
Another aspect of a human overseer that could potentially come to mind is their competence. Intuitively, a competent human overseer will be better at being effective. While this is true, there is good reason why competence is not part of the definition of effectiveness. Roughly, we could say that someone is competent in doing a task if they know what is the appropriate thing to do to fulfill the task, and are actually able to do it. So, competence can be reduced to the conditions of effectiveness, especially to epistemic access, self-control, and partly to causal power. Accordingly, a competent human overseer will be more likely to establish and retain these three conditions in a robust way, even though their competence is not part of their effectiveness per definition.
@@ 3.2.2. Responsibility
Our conceptualization of effectiveness in human oversight already accounts for responsibility while also going beyond that. Taking inspiration from a definition of moral responsibility in computing in (Noorman, 2020) and being consistent with the general idea of other sources (Talbert, 2019) we can define responsibility as follows: An agent is (morally) responsible for some aspect of the world if and only if (1) the agent has the power to establish a sufficient causal connection to this aspect of the world, (2) the agent has sufficient knowledge of their decision situation, and (3) the agent can decide for any path of action in their decision situation and, if they do so, follow through with it. In other words: a human overseer is morally responsible for risk-mitigation (or any other part of their job) if and only if they have (1) causal power, (2) epistemic access, and (3) self-control. This can be easily translated to cover past events: a human overseer was responsible for, say, mitigating a risk if and only if they had sufficient causal power, epistemic access, and self-control. These insights leads to the more concise formula that
A human overseer is effective if and only if they are morally responsible and have fitting intentions. [^6] If Definition 3.3 were adapted to say “could”, instead of “can” as suggested in footnote 5, this slogan would need to be altered to exclude cases of split-second, reflexive movements.
This account also allows for shared responsibility between multiple agents, which is desirable in the context of high-risk AI systems. If, for example, both a human overseer and the deployer had causal power, epistemic access, and self-control with regard to a certain risk, they would both be responsible. For reasons of simplicity, the definition does not allow for group responsibility, but can be adapted to do so.777 Our conceptualization of responsibility should not be considered a blueprint for the legal domain. While it may share similarities with some (though not all) existing legal definitions of responsibility, caution is warranted against equating the two, since legal definitions of responsibility vary significantly across diverse areas of law, legal systems, and legal cultures.
@@ 4. Facilitators and Inhibitors of Effective Human Oversight
Beyond providing a general conceptualization, our proposed definition of effectiveness in human oversight has a practically useful dimension. Specifically, its conditions can stimulate ideas about possible facilitators and inhibitors of effectiveness. Facilitators are expected to promote effective oversight, while inhibitors are expected to hinder it. We organize them into three categories: technical design features, individual factors of human overseers, and environmental circumstances (see, e.g., (Parker and Grote, 2020) for a similar scheme). The purpose of the discussion in this section is to provide readers with a more concrete understanding of possible factors regarding effective human oversight in accordance with our suggestions. The layout of Table 1 could serve as a template for thinking through possible facilitators and inhibitors, which can be useful to designers and deployers of AI-based systems. The following overview of facilitators and inhibitors is not meant to be exhaustive, and some of them may partially overlap each other. Also, they generally do not refer to a specific use case and some may not be applicable to every situation.
4.1. Technical design features
As “technical design features” we subsume facilitators and inhibitors whose influence on effective human oversight results from the design of the system. This list could also include aspects such as interface and interaction design (Norman, 1983), runtime monitoring systems (Baumeister et al., 2020), fault tolerance (Koren and Krishna, 2020), or automated anomaly detection systems (Nassif et al., 2021).
Intervention options. For an overseer to have causal power over the system, there must be options that allow the overseer to intervene with, control, overwrite, or undo system decisions and actions. For example, an intervention option might be a stop button that allows human overseers to stop system operation at any time. Other examples include options to take over manual control or to initiate emergency protocols that result in fail-safe modes.
System adaptability. It can be beneficial to oversight, if overseers can adapt the system’s processes or its interface (Evangelista Belo et al., 2022; Sauer and Chavaillaz, 2017). For example, the overseer could reduce the speed at which the system produces its outputs or adapt the system’s interface to display additional information. This increases the overseer’s epistemic access if the adaptations help him to acquire and process the information better.
Preselection of outputs to review. There may be cases where oversight tasks require a mechanism for selecting system outputs that are to be reviewed by the overseer. By defining boundary conditions or by using oversight support tools, human overseers could gain better epistemic access if only having to review a subset of outputs. For example, in a hiring context, such a tool could be used identify applicants for whom the AI-based evaluation was obviously fair and free of bias, leaving considerably fewer cases to the overseer (Biewer et al., 2024).
     intervention options      system adaptability      system understandability      interpretability of in- and outputs      preselection of outputs to review      overseer training      domain expertise      conscientiousness      exhaustion      motivation      automation bias      adequate job design      role conflicts      independent thinking      accountability      time pressure
@@ 4.1. Technical design features
4.2. Individual factors of human overseers
Under “individual factors of human overseers” we subsume traits and states of human overseers as well as interventions on human overseers that could facilitate or inhibit effective oversight. This discussion may cover aspects like vigilance (Warm et al., 2008), cognitive abilities (Spearman, 1904; Warne and Burningham, 2018), overseers’ propensity to trust in automation (Hoff and Bashir, 2015), and additional cognitive biases (Roese and Vohs, 2012) and heuristics (Tversky and Kahneman, 1973).
Overseer Training. Training to prepare for or enhance the role of the human overseer will be crucial. For instance, a training in which people are confronted with system errors or erroneous automated decisions can help them to become better at identifying these errors when later acting on the real system (Bahner et al., 2008). Overseer training could in principle target all of the conditions of effective human oversight. Such training could aim to enable overseers to maintain epistemic access, to recognize earlier when they are losing self-control (e.g., become inattentive (Perels et al., 2005)), to perform certain maneuvers that increase their causal power, or to know what the job requires them to do and thereby improve the fit of their intentions.
@@ 4.2. Individual factors of human overseers
Conscientiousness. Conscientiousness is a character trait of people who are self-disciplined, orderly, goal-directed, who make and follow plans, and who adhere to social norms (Costa et al., 1991). Highly conscientious overseers will be more likely to maintain self-control, sufficient epistemic access, and fitting intentions.
Exhaustion. Exhaustion is one of many psychological strain symptoms (Sonnentag et al., 2023). It is a state where an overseer feel tired and worn out (Maslach et al., 1997) and thus may not be good at maintaining self-control and sufficient epistemic access.
@@
Automation bias. Automation bias is usually considered a cognitive bias where people use the outcome of an automated decision aid “as a heuristic replacement for vigilant information seeking and processing” ((Mosier et al., 1996), p. 205). Automation bias becomes especially likely when facing a system that works reliably and with high performance because in such cases it is rational for people to divert their attention to other tasks (Parasuraman and Manzey, 2010). This, however, will contribute to insufficient epistemic access. If there are indeed no negative consequences for some time, this may lead to what Parasuraman and Manzey (Parasuraman and Manzey, 2010) called “learned carelessness”. This phenomenon could be interpreted as a problem of self-control or as an issue that leads to intentions no longer fitting the duties of a human overseer.

### v2

@@ 3.1.1. Causal power
Applied to the context of human oversight, this means that the oversight person has the power to establish a causal connection to parts of the system or its effects, especially the risk. For instance, this may include the ability to influence the system, including its components, operation, or outputs. Which kind of causal power is appropriate and necessary depends on the context at hand. For example, a factory robot assisting a human in assembly might need a stop button or a manual control mechanism. For an AI tool that assesses university applications and makes recommendations for the admission committee it might be enough if the oversight person can overwrite or disregard the system’s output and then forward possible issues to the system provider.
In any case, it is not enough if the oversight person is a mere observer of the system. Without proper means to interfere with it or at least with its effects, they could only watch the manifestation of a risk without being able to mitigate it. Also, it is not sufficient if the human only has ‘pseudo-options’, i.e., if they can only seemingly influence the system but not really make a relevant difference. [^3] Pseudo-options, though, are not to be confused with merely unattractive but real options, such as pushing a stop button for fun and thereby halting an entire production line for no reason. Instances of such scenarios are malfunctioning stop buttons, or environments in which the human’s actions are by default not considered.
@@ 3.1.2. Epistemic access
To be effective, the oversight person must know what to do and when to do it. More precisely:
@@
Applied to the context of human oversight, this means that the oversight person has sufficient knowledge about the risk and how to mitigate it, especially knowing a risk mitigation action. An oversight person will, for instance, need to have a sufficient understanding of what the system is doing, in which state it is, which possible risks and benefits there are, which means of influencing the system exist, which effects will result from possible interventions, and which of these effects are most desirable. In some cases, this might include knowledge of normative, social, or cultural domains, without which it would be infeasible for the oversight person to assess certain risks correctly. They do not need to have perfect knowledge, though. They might be unsure or even ignorant about some aspects of their decision situation – as long as they know enough to achieve their relevant objectives, they have enough epistemic access (Speith et al., ress). Notably, we assume that the oversight person has certain basic knowledge, such as what a stop button is or how to operate a computer on a consumer-level. One measure to improve the epistemic access of the oversight person could arguably be to employ suitable explainability and transparency measures (Sterz et al., 2021; Biewer et al., 2024; Schlicker et al., 2021; Baum et al., 2023; Jiang et al., 2022; Onitiu, 2023; Deck et al., 2024).
@@ 3.1.3. Self-control
To be effective, the oversight person must be in charge of their own doing. More precisely:
@@
Applied to the context of human oversight, this means that if there is a risk-mitigating path of action, the human can decide for that path, and if they do so, they actually perform the corresponding action(s). For this, the oversight person needs to be in their right state of mind and free in their actions in a relevant sense. They also have to be in a position to retain attention and purposefully target their actions during the oversight. If, for example, the job of overseeing a fleet of autonomous vehicles is so uneventful and dull that a normal person cannot help but let their mind wander off from their task, the oversight would ultimately become ineffective. Only if oversight persons have self-control, they can purposefully address risks. Other reasons why an oversight person could lack self-control would be when they are severely sleep-deprived, under the influence of substances, or suffering from an active seizure.
An oversight person who lacks self-control might also simultaneously lack sufficient causal power or epistemic access (e.g., in case of a seizure). Nevertheless, self-control is distinct from the other two conditions, as an oversight person can have self-control without having causal power or epistemic access. In fact, self-control may, in many situations, be an antecedent of causal power and epistemic access – but not vice versa.
Sometimes, oversight persons will decide in split-seconds or will instinctively perform a certain move, and we typically regard this as self-controlled. For example, an oversight person of a factory robot who realizes that it is just milliseconds away from severely injuring a human worker might reflexively slam the stop button. We consider this to be a real decision, even if made in a split-second; and since the oversight person followed suit with their decision, they are to be considered self-controlled. Therefore, even oversight persons deciding instinctively can be effective according to the upcoming Definition 3.5.555Alternatively, one might say there was not a true decision made in the example, but rather an instinctive jolt for the stop-button. In this case, the oversight person would not count as self-controlled according to Definition 3.3 (and therefore also not as effective according to Definition 3.5). If one wishes to allow this interpretation of the situation above and simultaneously allow that the human in it was exercising effective oversight, then the definition of self-control might be adapted to say “could” instead of “can”. In this case, even certain instinctive movements of the above kind fall under the definition. However, since we want to avoid a discussion of the conceptual idealization introduced by the “could”, we will stick to Definition 3.3 as is.
@@ 3.1.4. Fitting intentions
Applied to the context of human oversight, this means that the oversight person pro tanto intends to mitigate risks, i.e., they intend to mitigate any risks while taking other relevant factors, such as the interests of the system’s users and other stakeholders, into consideration in a suitable way. If the oversight person were unwilling to make the efforts necessary to mitigate a risk, they would not be effective in their oversight – even if they had causal power, epistemic access, and self-control. This may, for example, be the case if the oversight person lacks the motivation to do their job properly or has other conflicting interests. Even worse, if an oversight person had ill intent, they could use their position to foster risks instead of mitigating them. For example, a racist overseeing a university admission system could deliberately overwrite outputs where people of color are ranked highly. In this case, their oversight should be considered ineffective in the relevant respects.
@@ 3.1.5. Effectiveness
An oversight person is effective in their human oversight if and only if they have causal power, epistemic access, self-control, and fitting intentions in the above senses.
@@
Moreover, “effectiveness” is not a binary concept but a matter of degrees: one exercise of oversight might be more effective than another. Therefore, considering effectiveness in human oversight as a multifaceted continuum suitably encapsulates its nuanced nature. In line with the literature (Green, 2022) and legislation (European Commission, 2021), we nevertheless allow the use of “effective” as a binary concept. Saying that an oversight person is effective is to mean that they are effective enough for the concrete context they are operating in. In essence, this is also what Definition 3.5 says. While it is an open and important question of how much effectiveness is enough for a given context, we will not discuss this issue as it would go far beyond the scope of this paper.
@@ 3.2.1. Competence
Another aspect of an oversight person that could potentially come to mind is their competence. Intuitively, a competent oversight person will be better at being effective. While this is true, there is good reason why competence is not part of our definition of effectiveness: roughly speaking, someone is competent in performing a task if they know what to do to accomplish the task and are actually able to do it in a relevant sense. So, competence is tightly related to the conditions of effectiveness, even though competence is not itself part of the definition of effectiveness. A competent oversight person will be more likely to establish and retain sufficient epistemic access, self-control, and to use their causal powers in the right way. Without the necessary competence, an oversight person will arguably be unlikely to be effective, if they even can be effective at all.
@@ 3.2.2. Responsibility
Our conceptualization of effectiveness in human oversight already accounts for responsibility while also going beyond that. Taking inspiration from a definition of moral responsibility in computing in (Noorman, 2020) and being consistent with the general idea of other sources (Talbert, 2019; Fischer and Ravizza, 1998) we can define responsibility as follows: An agent is (morally) responsible for some aspect of the world if and only if (1) the agent has the power to establish a sufficient causal connection to this aspect of the world, (2) the agent has sufficient knowledge of their decision situation, and (3) the agent can decide for any path of action in their decision situation and, if they do so, follow through with it. In other words: an oversight person is morally responsible for risk mitigation (or any other part of their job) if and only if they have (1) causal power, (2) epistemic access, and (3) self-control in that respect. This can be easily translated to cover past events: an oversight person was responsible for, say, mitigating risk if and only if they had sufficient causal power, epistemic access, and self-control in that respect. These insights lead to the more concise formula that
An oversight person is effective if and only if they are morally responsible and have fitting intentions. [^6] If Definition 3.3 were adapted to say “could”, instead of “can” as suggested in footnote 5, this slogan would need to be altered to exclude cases of split-second, reflexive movements.
This account also allows for shared responsibility between multiple agents, which is desirable in the context of high-risk AI systems. If, for example, both an oversight person and the deployer had causal power, epistemic access, and self-control with regard to a certain risk, they would both be responsible. For reasons of simplicity, the definition does not allow for group responsibility but can be adapted to do so.777 Our conceptualization of responsibility should not be considered a blueprint for the legal domain. While it may share similarities with some (though not all) existing legal definitions of responsibility, caution is warranted against equating the two, since legal definitions of responsibility vary significantly across diverse areas of law, legal systems, and legal cultures.
@@ 4. Facilitators and Inhibitors of Effective Human Oversight
Beyond providing a general conceptualization, our proposed definition of effectiveness in human oversight has a practically useful dimension. Specifically, its conditions can stimulate ideas about possible facilitators and inhibitors of effectiveness. Facilitators are expected to promote effective oversight, while inhibitors are expected to hinder it. We organize them into three categories: technical design features, individual factors of oversight persons, and environmental circumstances (see, e.g., (Parker and Grote, 2020) for a similar scheme). The purpose of the discussion in this section is to provide readers with a more concrete understanding of possible factors regarding effective human oversight in accordance with our suggestions. The layout of Table 1 could serve as a template for thinking through possible facilitators and inhibitors, which can be useful to designers and deployers of AI-based systems. The following overview of facilitators and inhibitors is not meant to be exhaustive, and some of them may partially overlap each other. Also, they generally do not refer to a specific use case and some may not be applicable to every situation.
intervention options
system adaptability
system understandability
interpretability of in- and outputs
preselection of outputs to review
training of the oversight person
domain expertise
conscientiousness
exhaustion
motivation
automation bias
adequate job design
role conflicts
independent thinking
accountability
time pressure
@@
4.1. Technical design features
Under “technical design features”, we subsume facilitators and inhibitors whose influence on effective human oversight results from the design of the system. This list could also include aspects such as interface and interaction design (Norman, 1983), runtime monitoring systems (Baumeister et al., 2020), fault tolerance (Koren and Krishna, 2020), or automated anomaly detection systems (Nassif et al., 2021).
Intervention options. For an oversight person to have causal power over the system, there must be options that allow the oversight person to intervene with, control, overwrite, or undo system decisions and actions. For example, an intervention option might be a stop button that allows oversight persons to stop system operation at any time. Other examples include options to take over manual control or to initiate emergency protocols that result in fail-safe modes.
System adaptability. It can be beneficial to oversight if oversight persons can adapt the system’s processes or its interface (Evangelista Belo et al., 2022; Sauer and Chavaillaz, 2017). For example, the oversight person could reduce the speed at which the system produces its outputs or adapt the system’s interface to display additional information. This increases the oversight person’s epistemic access if the adaptations help him to acquire and process the information better.
Preselection of outputs to review. There may be cases where oversight tasks require a mechanism for selecting system outputs that are to be reviewed by the oversight person. By defining boundary conditions or by using oversight support tools, oversight persons could gain better epistemic access when only having to review a subset of outputs. For example, in a hiring context, such a tool could be used to identify applicants for whom the AI-based evaluation was obviously fair and free of bias, leaving considerably fewer cases to the oversight person (Biewer et al., 2024).
4.2. Individual factors of oversight persons
Under “individual factors of oversight persons” we subsume traits and states of oversight persons as well as interventions on oversight persons that could facilitate or inhibit effective oversight. This discussion may cover aspects like vigilance (Warm et al., 2008), cognitive abilities (Spearman, 1904; Warne and Burningham, 2018), humans’ propensity to trust in automation (Hoff and Bashir, 2015), and additional cognitive biases (Roese and Vohs, 2012) and heuristics (Tversky and Kahneman, 1973).
Overseer Training. Training to prepare for or enhance the role of the oversight person will be crucial. For instance, training in which people are confronted with system errors or erroneous automated decisions can help them to become better at identifying these errors when later acting on the real system (Bahner et al., 2008). Overseer training could, in principle, target all of the conditions of effective human oversight. Such training could aim to enable oversight persons to maintain epistemic access, to recognize earlier when they are losing self-control (e.g., become inattentive (Perels et al., 2005)), to perform certain maneuvers that increase their causal power, or to know what the job requires them to do and thereby improve the fit of their intentions.
@@ 4.2. Individual factors of oversight persons
Conscientiousness. Conscientiousness is a character trait of people who are self-disciplined, orderly, goal-directed, who make and follow plans, and who adhere to social norms (Costa et al., 1991). Highly conscientious oversight persons will arguably be more likely to maintain self-control, sufficient epistemic access, and fitting intentions.
Exhaustion. Exhaustion is one of many psychological strain symptoms (Sonnentag et al., 2023). It is a state where an oversight person feels tired and worn out (Maslach et al., 1997) and thus may not be good at maintaining self-control and sufficient epistemic access.
@@
Automation bias. Automation bias is usually considered a cognitive bias where people use the outcome of an automated decision aid “as a heuristic replacement for vigilant information seeking and processing” ((Mosier et al., 1996), p. 205). Automation bias becomes especially likely when facing a system that works reliably and with high performance because, in such cases, it is rational for people to divert their attention to other tasks (Parasuraman and Manzey, 2010). This, however, will contribute to insufficient epistemic access. If there are indeed no negative consequences for some time, this may lead to what Parasuraman and Manzey (Parasuraman and Manzey, 2010) called “learned carelessness”. This phenomenon could be interpreted as a problem of self-control or as an issue that leads to intentions that no longer fit the duties of an oversight person.